Find the measure of an angle whose supplement measures 5 degrees greater than 6 times its complement. If I could figure out the correct formula I would be OK. Thanks.
Answer by Earlsdon(6294)

Let the angle = A.
Its supplement is (180-A) and its complement is (90-A).
From the problem description you can write"
(180-A) = 6(90-A)+5 Simplify and solve for A
180-A = 540-6A+5
180-A = 545-6A Add 6A to both sides.
180+5A = 545 Subtract 180 from both sides.
5A = 365 Divide both sides by 5.
A = 73 degrees.